OpenStudy (anonymous):

What is equal to square root of the cube root of 6?

OpenStudy (jdoe0001):

\(\large \bf \sqrt{\sqrt[3]{6}}\quad ?\)

OpenStudy (solomonzelman):

\[\sqrt{\sqrt[3]{6}}=(6^{1/2})^{1/3}\]

OpenStudy (anonymous):

thts the answer?

OpenStudy (solomonzelman):

no, it's no simplified.

OpenStudy (jdoe0001):

\(\huge \bf \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)

OpenStudy (anonymous):

@SolomonZelman thats how i can slove it?

OpenStudy (solomonzelman):

what is how you solved it? I don't see any work posted by you

OpenStudy (jdoe0001):

recall your exponent rules \(\large \bf (x^a)^b\implies x^{a\cdot b}\)

OpenStudy (anonymous):

@jdoe0001 oooh i see it now :)